# Homological properties of the relative Frobenius morphism

### v1

@@ Abstract.
This work concerns maps of commutative noetherian local rings containing a field of positive characteristic. Given such a map $\varphi$ of finite flat dimension, the results relate homological properties of the relative Frobenius of $\varphi$ to those of the fibers of $\varphi$. The focus is on the complete intersection property and the Gorenstein property.
@@ 1. Introduction
[…]
@@
Assuming $R$ is a local ring with residue field $k$, the derived fiber of $\varphi$ refers to $S\otimes_{R}^{\mathbb{L}}k$, thought of as a simplicial $k$-algebra. When $\varphi$ is flat, this is simply $S\otimes_{R}k$. The Frobenius extends naturally to the setting of simplicial rings by applying the classical Frobenius degreewise, and so we can consider $F\colon S\otimes_{R}^{\mathbb{L}}k\to S\otimes_{R}^{\mathbb{L}}k$. This leads us to our first result.
Theorem 1.1.
Let $R$ and $S$ be $F$-finite local rings of positive characteristic. If $\varphi\colon R\to S$ is a local homomorphism of finite flat dimension, the Betti numbers of the relative Frobenius $F_{S/R}$ grow at the same rate as the Betti numbers of the Frobenius on $S\otimes_{R}^{L}k$.
A precise formulation of this result can be found in Theorem 3.1. As a corollary, we recover the results of Radu [Rad92], André [And93, And94] and Dumitrescu [Dum96] (concerning regularity) and Blanco-Majadas [BM98] (concering the complete intersection property) while also weakening the hypothesis that $\varphi$ is flat to requiring that $\varphi$ has finite flat dimension.
Corollary 1.2.
For $\varphi$ as in Theorem 3.1, the fibers of $\varphi$ are regular (resp. complete intersection) if and only if $F_{S/R}$ has finite flat (resp. CI-) dimension.
We also prove a similar result for the Gorenstein property.
Theorem 1.3.
Let $R$ and $S$ be $F$-finite local rings of positive characteristic. If $\varphi\colon R\to S$ is a flat local homomorphism, then the fibers of $\varphi$ are Gorenstein if and only if $F_{S/R}$ has finite G-dimension.
Note that we require $\varphi$ to be flat here rather than finite flat dimension as in the previous results. The reason for this is that for $\varphi$ of finite flat dimension, we must consider the derived fibers of $\varphi$. We plan to take this up on a later occasion.
Acknowledgements
This material is based on work supported by the National Science Foundation under grants DMS-200985 and RTG DMS-1840190. The author would like to thank his advisor Srikanth Iyengar for many helpful comments and conversations and his continued support in the writing of this paper, as well as Rankeya Datta, Linquan Ma, Daniel McCormick, Brendan Murphy, and Karl Schwede for additional helpful conversations.
@@ 2. Simplicial rings
that preserves the natural (projective) model structures, that is to say, it is a Quillen equivalence. If we upgrade the setting to a simplicial ring $A$, then $N(A)$ is a differential graded ring and we can consider $\operatorname{Mod}_{\geq 0}(N(A))$, the category of dg $N(A)$-modules in nonnegative homological degree. [SS03, Theorem 1.1] gives a Quillen equivalence between $\operatorname{Mod}(A)$ and $\operatorname{Mod}_{\geq 0}(N(A))$ where $N$ is the right adjoint (note that $\Gamma$ is not a left adjoint in this case [SS03, Section 3.3]). Homotopy groups of an $A$-module can be computed by first applying the normalization functor and computing homology as a dg $N(A)$-module, so we will use these equivalences to move freely between the simplicial and dg settings.
@@
Let $A$ be a commutative simplicial ring. We say that $A$ is local if $A_{0}$ is a noetherian local ring. If $k$ is the residue field of $A_{0}$, we write $(A,k)$. Given $M\in\operatorname{Mod}(A)$, we say that $M$ is a finite $A$-module if $\pi_{i}(M)$ is a finite $\pi_{0}(A)$-module for each $i$ and $\pi_{i}(M)=0$ for $i\gg 0$. We write $\operatorname{mod}(A)$ for the category of finite $A$-modules.
@@
$\operatorname{curv}_{A}(M\otimes_{B}^{\mathbb{L}}N)\leq\max\{\operatorname{curv}_{A}(M),\operatorname{curv}_{B}(N)\}$
@@
$E^{2}_{p,q}\colon=\pi_{p}(N\otimes_{B}^{\mathbb{L}}\pi_{q}(M\otimes_{A}^{\mathbb{L}}k_{A}))\implies\pi_{p+q}((M\otimes_{B}^{\mathbb{L}}N)\otimes_{A}^{\mathbb{L}}k_{A})$
@@
$\pi_{p}(N\otimes_{B}^{\mathbb{L}}\pi_{q}(N\otimes_{A}^{\mathbb{L}}k_{A}))\cong\pi_{p}(N\otimes_{B}^{\mathbb{L}}k_{B})\otimes_{k_{A}}\pi_{q}(M\otimes_{A}^{\mathbb{L}}k_{A}))$
@@
Let $(A,k_{A})$ and $(B,k_{B})$ be local simplicial rings and let $M\in\operatorname{mod}(B)$. If $\varphi\colon A\to B$ is a flat map of local simplicial rings then
@@
In this section, we collect some results about simplicial rings in positive characteristic that will help contextualize our later results on homological dimension. We first prove the following result which is simply a re-framing of [${}^{+}$23, Theorem 2.1] for simplicial rings. The proof is exactly the same aside from checking that various statements remain true when we replace a classical ring $R$ with a simplicial ring $A$. We include it here for the convenience of the reader.
@@
From this we can derive a simplicial version of Theorem 1.1 of [AHIY12], itself a generalization of [Kun69, Theorem 2.1] and [Rod88, Theorem 2].
Proposition 2.12.
Let $(A,k)$ be a local simplicial ring and $M\in\operatorname{mod}(A)$. If $\operatorname{curv}_{A}(F^{e}_{*}M)=0$, then $A$ has finite global dimension and thus $A\simeq\pi_{0}(A)$ is a regular ring.
Proof.
If $\operatorname{curv}_{A}(F^{e}_{*}M)$, then $\operatorname{curv}_{N(A)}N(F^{e}_{*}M)=0$ by Lemma 2.4. Then $N(F^{e}_{*}M)$ and $N(F^{e}_{*}K^{M})$ are in $\operatorname{thick}(N(A))$. Then, by Proposition 2.11, $k\in\operatorname{thick}(N(A))$. Then, by [Jør10, Theorem A], $A\simeq\pi_{0}(A)$ and so $\pi_{0}(A)$ is regular. ∎
We can also use Proposition 2.11 to extend, with similar proof, [AHIY12, Theorem 5.1].
@@ 3. Homological dimension
This section contains the main results connecting the homological properties of the relative Frobenius of a map $\varphi$ of local rings to the homological properties of the (derived) fibers of $\varphi$. We first focus on regularity and the complete intersection property, as these are characterized in terms of Betti number growth, before turning to the Gorenstein property.
Growth of Betti numbers
Throughout, all rings will be commutative, noetherian of characteristic $p>0$. We will also assume all rings are $F$-finite. Fix a local homomorphism $\varphi\colon(R,\mathfrak{m}_{R},k_{R})\to(S,\mathfrak{m}_{S},k_{S})$ of finite flat dimension. In this setting, a result of Peskine-Szpiro [PS73] gives that $S\otimes_{R}^{\mathbb{L}}F_{*}R\cong S\otimes_{R}F_{*}R$. Our goal is to understand how the homological properties of the relative Frobenius relate to those of the (derived) fibers of $\varphi$. Recall the following diagram
[…]
@@
Let $R$ and $S$ be $F$-finite local rings of positive characteristic with residue fields $k_{R}$ and $k_{S}$ respectively. Let $\varphi\colon R\to S$ be a map of local rings. If $\varphi$ has finite flat dimension, then
@@
where $\bar{S^{\prime}}:=S\otimes_{R}^{\mathbb{L}}k^{\prime}$ for $k^{\prime}$ any finite, purely inseparable extension of $k_{R}$ and $A$ is the localization of $S\otimes_{R}F_{*}R$ at the contraction of the maximal ideal of $F_{*}S$.
@@
Let $k^{\prime}$ be a finite, purely inseparable field extension of $k_{R}$ and let $e$ be such that $k^{\prime}\subseteq F^{e}_{*}k_{R}$. Let $A^{e}$ be the localization of $S\otimes_{R}F^{e}_{*}R$ at the contraction of the maximal ideal of $F^{e}_{*}S$, let $k_{A^{e}}$ be the residue field of $A^{e}$, and fix the following notation:
@@
Suppose $F_{S/R}$ has finite flat dimension. Regularity is a local condition, so we can work in the setting of Theorem 3.1 and assume $R$ and $S$ are local. Let $k^{\prime}$ be a finite, purely inseparable field extension of $k_{R}$, the residue field of $R$ and let $\bar{S^{\prime}}:=S\otimes_{R}^{\mathbb{L}}k^{\prime}$. Because $\beta_{i}^{A}(F_{*}S)=0$ for $i\gg 0$, Theorem 3.1 tells us $\beta_{i}^{\bar{S}}(F_{*}\bar{S})=0$ for $i\gg 0$. By Proposition 2.12, $\bar{S^{\prime}}\cong S\otimes_{R}k^{\prime}$ is regular and flatness of $\varphi$ follows immediately. The reverse direction follows from [Rad92, And93]. ∎
@@
We say that a map of local rings $\varphi\colon R\to S$ is complete intersection at the maximal ideal of $S$ if for some (equiv. any) Cohen factorization $R\to R^{\prime}\to\hat{S}$ of $\varphi$, the ideal $\ker(R^{\prime}\to\hat{S})$ is generated by a regular sequence. $\mathbf{f^{\prime}}$.
@@
    the differential of A satisfies $\partial(\mathfrak{m}A)\subseteq\mathfrak{m}_{R}A+\mathfrak{m}_{A}^{2}$;
@@
Let $R$ and $S$ be $F$-finite local rings of positive characteristic. Let $\varphi\colon R\to S$ be a map of local rings of finite flat dimension. Then $F_{S/R}$ has finite CI dimension if and only if $\varphi$ is complete intersection at $\mathfrak{m}_{S}$.
@@
Let $A$ be the localization of $S\otimes_{R}F_{*}S$ at the contraction of the maximal ideal of $F_{*}S$. Let $\bar{S}\colon=S\otimes_{R}^{\mathbb{L}}k_{R}$ where $k_{R}$ is the residue field of $R$. Since $F_{S/R}$ has finite CI-dimension, $\operatorname{curv}_{A}(F_{*}S)\leq 1$, by [AGP97, Theorem 5.3], and hence $\operatorname{curv}_{\bar{S}}(k_{S})\leq 1$, by Theorem 3.1. We claim that this implies $\varphi$ is complete intersection at $\mathfrak{m}_{S}$; to deduce this we use the results from [Avr99, Theorem 3.4]; see also [AI03, Theorem 5.4].
@@
[…]
@@
G-dimension
With a Radu-André-type result for flat and CI-dimension, it is natural to ask whether a similar result can be proven for G-dimension, an analogue of CI-dimension for the Gorenstein property. This was first defined for finite modules in [AB69] and was generalized to modules over local homomorphisms in [ISW04].
Definition 3.9.
Given $M$ a homologically finite complex of $R$-modules, meaning $H(M)$ is degreewise finite and bounded, we say $M$ has finite $G$-dimension if the following natural map is an isomorphism
[…]
That is to say, $M$ is derived reflexive.
In the following result, we restrict to a flat homomorphism $\varphi\colon R\to S$, rather than requiring $\varphi$ to have finite flat dimension as in the previous section. This is to avoid discussing the definition of finite $G$-dimension in the setting of derived rings which we plan to take up on a later occasion.
Theorem 3.10.
Let $R$ and $S$ be $F$-finite local rings of positive characteristic. Let $\varphi\colon R\to S$ be a flat map of local rings. Then the relative Frobenius $F_{S/R}\colon S\otimes_{R}F_{*}R\to F_{*}S$ has finite G-dimension if and only if $\varphi$ is Gorenstein at $\mathfrak{m}_{S}$.
Proof.
Recall the following diagram
[…]
Since the $(1)$ is just the base change of a field extension, if we can show the $(2)$ has finite $G$-dimension, we will be done. So, we simply want to show that finite $G$-dimension is preserved when we base change along $F_{*}R\to F_{*}k_{R}$. Let $A\colon=S\otimes_{R}F_{*}R$. Since the relative Frobenius has finite $G$-dimension, we consider the following isomorphism
[…]
Applying $-\otimes^{\mathbb{L}}_{F_{*}R}F_{*}k_{R}$, we get
[…]
Then, because $\varphi$ is flat,
[…]
where again $\bar{A}\colon=S\otimes_{R}F_{*}k_{R}$.
For the reverse direction, consider the exact triangle in $\operatorname{Ho}(R)$
$F_{*}S\to\operatorname{RHom}_{A}(\operatorname{RHom}_{A}(F_{*}S,A),A)\to C$
where $C$ is the cone of the natural map $F_{*}S\to\operatorname{RHom}_{A}(\operatorname{RHom}_{A}(F_{*}S,A),A)$. If, after applying $-\otimes_{R}^{\mathbb{L}}F_{*}k_{R}$ to the triangle, we get that the first map is a quasi-isomorphism, this implies $C\otimes_{R}^{\mathbb{L}}F_{*}k_{R}$ is acyclic. By Nakayama’s Lemma, this implies $C$ is acyclic.
Thus we have that $S_{S/R}$ has finite $G$-dimension if and only if $F:S\otimes_{R}k_{R}\to F_{*}(S\otimes_{R}k_{R})$ has finite $G$-dimension, and the result follows from [TY04, Theorem 6.2]. ∎

### v2

@@ Abstract.
This work concerns maps of commutative noetherian local rings containing a field of positive characteristic. Given such a map $\varphi$ of finite flat dimension, the results relate homological properties of the relative Frobenius of $\varphi$ to those of the fibers of $\varphi$ with a focus on the complete intersection property.
@@ 1. Introduction
${R}$${S}$${F_{*}R}$${S\otimes_{R}F_{*}R}$${F_{*}S}$$\scriptstyle{\varphi}$$\scriptstyle{F}$$\scriptstyle{F}$$\scriptstyle{F_{S/R}}$
@@
Assuming $R$ is a local ring with residue field $k$, the derived fiber of $\varphi$ refers to $S\otimes_{R}^{\mathbb{L}}k$, thought of as a simplicial $k$-algebra. When $\varphi$ is flat, this is simply $S\otimes_{R}k$. The Frobenius extends naturally to the setting of simplicial rings by applying the classical Frobenius degreewise, and so we can consider $F\colon S\otimes_{R}^{\mathbb{L}}k\to S\otimes_{R}^{\mathbb{L}}k$. This leads us to our main result.
Theorem 1.1 (3.1).
Let $\varphi\colon R\to S$ be a map of $F$-finite local rings of positive characteristic and let $k_{R}$ be the residue field of $R$. If $\varphi$ has finite flat dimension, then
where $\bar{S^{\prime}}:=S\otimes_{R}^{\mathbb{L}}k^{\prime}$ for $k^{\prime}$ any finite, purely inseparable extension of $k_{R}$.
Here, curvature of a finitely generated module $M$ over a local ring $T$ is defined as
$\operatorname{curv}_{T}(M):=\limsup_{n}\sqrt[n]{\beta^{T}_{n}(M)}$
and the definition can be extended to modules over simplicial rings (see Definitions 2.2 and 2.3). Curvature measures the exponential growth rate of the Betti numbers of $M$, so in more informal language this result shows that the Betti numbers of the relative Frobenius $F_{S/R}$ grow at the same rate as the Betti numbers of the Frobenius on $S\otimes_{R}^{\mathbb{L}}k$.
As a corollary, we recover the results of Radu [Rad92], André [And93, And94] and Dumitrescu [Dum96] (concerning regularity) and Blanco-Majadas [BM98] (concering the complete intersection property) while also weakening the hypothesis that $\varphi$ is flat to requiring that $\varphi$ has finite flat dimension.
Corollary 1.2 (3.3, 3.8).
Let $\varphi\colon R\to S$ be a map of $F$-finite local rings of positive characteristic and $\mathfrak{m}_{S}$ be the maximal ideal of $S$. If $\varphi$ has finite flat dimension, then $\varphi$ is regular (resp. complete intersection at the maximal ideal of $S$) if and only if $F_{*}S$ has finite flat (resp. CI-) dimension over $S\otimes_{R}F_{*}R$.
@@ 2. Simplicial rings
that preserves the natural (projective) model structures, that is to say, it is a Quillen equivalence. If we upgrade the setting to a simplicial ring $A$, then $N(A)$ is a differential graded ring and we can consider $\operatorname{Mod}_{\geq 0}(N(A))$, the category of dg $N(A)$-modules in nonnegative homological degree. [SS03, Theorem 1.1] gives a Quillen equivalence between $\operatorname{Mod}(A)$ and $\operatorname{Mod}_{\geq 0}(N(A))$ where $N$ is the right adjoint (note that $\Gamma$ is not a left adjoint in this case [SS03, Section 3.3]). Homotopy groups $\pi_{*}(M)$ of an $A$-module $M$ can be computed by first applying the normalization functor and computing homology of $N(M)$ as a dg $N(A)$-module, so we will use these equivalences to move freely between the simplicial and dg settings.
@@
Let $A$ be a commutative simplicial ring. We say that $A$ is local if $A_{0}$ is a noetherian local ring, in which case $\pi_{0}(A)$ is also a noetherian local ring. If $k$ is the residue field of $A_{0}$, we write $(A,k)$. Given $M\in\operatorname{Mod}(A)$, we say that $M$ is a finite $A$-module if $\pi_{i}(M)$ is a finite $\pi_{0}(A)$-module for each $i$ and $\pi_{i}(M)=0$ for $i\gg 0$. We write $\operatorname{mod}(A)$ for the category of finite $A$-modules.
@@
$\operatorname{curv}_{A}(M\otimes_{B}^{\mathbb{L}}N)\leq\max\{\operatorname{curv}_{A}(M),\operatorname{curv}_{B}(N)\}.$
@@
$E^{2}_{p,q}\colon=\pi_{p}(N\otimes_{B}^{\mathbb{L}}\pi_{q}(M\otimes_{A}^{\mathbb{L}}k_{A}))\implies\pi_{p+q}((M\otimes_{B}^{\mathbb{L}}N)\otimes_{A}^{\mathbb{L}}k_{A}).$
@@
$\pi_{p}(N\otimes_{B}^{\mathbb{L}}\pi_{q}(M\otimes_{A}^{\mathbb{L}}k_{A}))\cong\pi_{p}(N\otimes_{B}^{\mathbb{L}}k_{B})\otimes_{k_{A}}\pi_{q}(M\otimes_{A}^{\mathbb{L}}k_{A})).$
@@
Let $(A,k_{A})$ and $(B,k_{B})$ be local simplicial rings and let $M\in\operatorname{mod}(B)$. If $\varphi\colon A\to B$ is a finite flat map of local simplicial rings then
@@
In this section, we collect some results about simplicial rings in positive characteristic that will help contextualize our later results on homological dimension. We first prove the following result which is simply a re-framing of [BIL+23, Theorem 2.1] for simplicial rings. The proof is exactly the same aside from checking that various statements remain true when we replace a classical ring $R$ with a simplicial ring $A$. We include it here for the convenience of the reader.
@@
It’s worth noting that the $c$ in the statement of Proposition 2.11 is an invariant of $A$ and thus does not depend on the module $M$. We can use Proposition 2.11 to extend, with similar proof, [AHIY12, Theorem 5.1].
@@
From this we can also derive a simplicial version of Theorem 1.1 of [AHIY12], itself a generalization of [Kun69, Theorem 2.1] and [Rod88, Theorem 2].
Proposition 2.13.
Let $(A,k)$ be a local simplicial ring and $M\in\operatorname{mod}(A)$. If $\operatorname{curv}_{A}(F^{e}_{*}M)=0$ for some (equiv. for all) $e\geq 1$, then $A$ has finite global dimension and thus $A\simeq\pi_{0}(A)$ is a regular ring.
Proof.
By Lemma 2.12, $\operatorname{curv}_{A}(F^{e}_{*}M)=\operatorname{curv}_{A}(k)$ for all $e\geq 1$ and so $\operatorname{curv}_{A}(k)=0$. Then $\operatorname{curv}_{N(A)}(k)=0$ by Lemma 2.4 so $k\in\operatorname{thick}(N(A))$. Then, by [Jør10, Theorem A], $A\simeq\pi_{0}(A)$ and so $\pi_{0}(A)$ is regular. ∎
@@ 3. Homological dimension
This section contains the main results connecting the homological properties of the relative Frobenius of a map $\varphi$ of local rings to the homological properties of the (derived) fibers of $\varphi$. We focus on regularity and the complete intersection property, as these are characterized in terms of Betti number growth.
Throughout, all rings will be commutative, noetherian of characteristic $p>0$. We will also assume all rings are $F$-finite. Fix a local homomorphism $\varphi\colon(R,\mathfrak{m}_{R},k_{R})\to(S,\mathfrak{m}_{S},k_{S})$ of finite flat dimension. In this setting, [MW16, Corollary 3.5(a)] gives that $S\otimes_{R}^{\mathbb{L}}F_{*}R\cong S\otimes_{R}F_{*}R$ when $R\to S$ is finite. Our goal is to understand how the homological properties of the relative Frobenius relate to those of the (derived) fibers of $\varphi$. Recall the following diagram
${R}$${S}$${F_{*}R}$${A\colon=S\otimes_{R}F_{*}R}$${F_{*}S}$$\scriptstyle{\varphi}$$\scriptstyle{F}$$\scriptstyle{F}$$\scriptstyle{F_{S/R}}$
@@
Let $\varphi\colon R\to S$ be a map of $F$-finite local rings of positive characteristic and let $k_{R}$ be the residue field of $R$. If $\varphi$ has finite flat dimension, then
@@
where $\bar{S^{\prime}}:=S\otimes_{R}^{\mathbb{L}}k^{\prime}$ for $k^{\prime}$ any finite, purely inseparable extension of $k_{R}$.
@@
Let $k^{\prime}$ be a finite, purely inseparable field extension of $k_{R}$ and let $e$ be such that $k^{\prime}\subseteq F^{e}_{*}k_{R}$. Let $A^{e}\colon=S\otimes_{R}F^{e}_{*}R$. Note that by [Gro77, Exp. XV Propostion 2a], the relative Frobenius induces a homeomorphism on spectra so $A^{e}$ is local. Let let $k_{A^{e}}$ be the residue field of $A^{e}$. We fix the following notation:
@@
Suppose $F_{S/R}$ has finite flat dimension. Let $k^{\prime}$ be a finite, purely inseparable field extension of $k_{R}$, the residue field of $R$ and let $\bar{S^{\prime}}:=S\otimes_{R}^{\mathbb{L}}k^{\prime}$. Because $\beta_{i}^{A}(F_{*}S)=0$ for $i\gg 0$, Theorem 3.1 tells us $\beta_{i}^{\bar{S}}(F_{*}\bar{S})=0$ for $i\gg 0$. By Proposition 2.13, $\bar{S^{\prime}}\cong S\otimes_{R}k^{\prime}$ is regular and flatness of $\varphi$ follows immediately. The reverse direction follows from [Rad92, And93]. ∎
@@
We say that a map of local rings $\varphi\colon R\to S$ is complete intersection at the maximal ideal of $S$ if for some (equiv. any) Cohen factorization $R\to R^{\prime}\to\hat{S}$ of $\varphi$, the ideal $\ker(R^{\prime}\to\hat{S})$ is generated by a regular sequence $\mathbf{f^{\prime}}$.
@@
    the differential of A satisfies $\partial(\mathfrak{m}_{R}A)\subseteq\mathfrak{m}_{R}A+\mathfrak{m}_{A}^{2}$;
@@
Let $\varphi\colon R\to S$ be a map of $F$-finite local rings of positive characteristic and let $\mathfrak{m}_{S}$ be the maximal ideal of $S$. If $\varphi$ has finite flat dimension, then $F_{*}S$ has finite CI dimension over $S\otimes_{R}F_{*}R$ if and only if $\varphi$ is complete intersection at $\mathfrak{m}_{S}$.
@@
Let $A\colon=S\otimes_{R}F_{*}S$ and let $\bar{S}\colon=S\otimes_{R}^{\mathbb{L}}k_{R}$ where $k_{R}$ is the residue field of $R$. Since $F_{S/R}$ has finite CI-dimension, $\operatorname{curv}_{A}(F_{*}S)\leq 1$ by [AGP97, Theorem 5.3], and hence $\operatorname{curv}_{\bar{S}}(k_{S})\leq\operatorname{curv}_{\bar{S}}(F_{*}\bar{S})\leq 1$ by a combination of Theorem 3.1 and Lemma 2.12. We claim that this implies $\varphi$ is complete intersection at $\mathfrak{m}_{S}$; to deduce this we use the results from [Avr99, Theorem 3.4]; see also [AI03, Theorem 5.4].
@@
${R}$${S}$${R^{\prime}}$$\scriptstyle{\dot{\varphi}}$$\scriptstyle{\grave{\varphi}}$$\scriptstyle{\varphi^{\prime}}$